OpenStudy (anonymous):

How to solve the inequality:2x-1<=[x+1]

OpenStudy (jiteshmeghwal9):

2x-1< x+1 =2x-1+1<x+1+1 =2x<2+x =2x-x<2+x-x =x<2

OpenStudy (jiteshmeghwal9):

am I correct @mukushla

OpenStudy (anonymous):

i dont know what he/she means by [x+1] maybe its greatest integer less than or equal to x+1 !!!!

OpenStudy (anonymous):

if not u r right

OpenStudy (jiteshmeghwal9):

he/she only wants the solution of the inequality.